OpenStudy (anonymous):

Solve the equation of 2log (2y + 4) = log 9 + 2log (y -1). A. 3 B. 5 C. 7 D. 21

OpenStudy (anonymous):

@Hero @ParthKohli @iambatman

OpenStudy (anonymous):

first write both sides of the equation as a single log then ignore the logs

OpenStudy (anonymous):

that is one way, there may be an easier way

OpenStudy (anonymous):

Ok so then its B correct?

OpenStudy (anonymous):

@TheSmartOne @AlexandervonHumboldt2

OpenStudy (anonymous):

\[2\log(2y+4)=\log9+2\log(y-1)\]

OpenStudy (anonymous):

\[2\log(2y+4)-2\log(y-1)=\log9\]

OpenStudy (anonymous):

\[\log(2y+4)^2-\log(y-1)^2=\log9\]

OpenStudy (anonymous):

\[\log \frac{ (2y+4)^2 }{ (y-1)^2 }=\log9\]

OpenStudy (anonymous):

\[\log(\frac{ 2y+4 }{ y-1 })^2=\log9\]

OpenStudy (anonymous):

Then,cancel both log

OpenStudy (anonymous):

\[(\frac{ 2y+4 }{ y-1 })^2=9\]

OpenStudy (anonymous):

\[\frac{ 2y+4 }{ y-1 }=\sqrt{9}\]

OpenStudy (anonymous):

\[\frac{ 2y+4 }{ y-1 }=3\]

OpenStudy (anonymous):

Multiply both sides by y-1

OpenStudy (anonymous):

\[\frac{ 2y+4 }{ y-1 }\times~(y-1)=3\times(y-1)\]

OpenStudy (anonymous):

\[2y+4=3(y-1)\]

OpenStudy (anonymous):

Expand the right side

OpenStudy (anonymous):

\[2y+4=3y-3\]

OpenStudy (anonymous):

Then,combine like terms and solve for y...

OpenStudy (anonymous):

@brooke_ivy99